OpenStudy (anonymous):

6 + 10 x + 14 x^2 what would be the greatest common factor and is it smaller than the smallest coefficient

OpenStudy (anonymous):

gcf is 2 which is less than 6

OpenStudy (anonymous):

To find the GCF, factor each individual number into it's prime factors first. 6 = 2*3 10 = 2*5 14 = 2*7 The biggest factor that all these numbers have in common is 2. Therefore the GCF is 2.

OpenStudy (anonymous):

thats what i thought thanks all good night